OpenStudy (madeleine2):

Write the standard form of the line that passes through the given points (1, 1) and (3, 4)

OpenStudy (3mar):

May I help?

OpenStudy (madeleine2):

sure

563blackghost (563blackghost):

Find the slope of the points. \(\huge\bf{\frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{4-1}{3-1}=\color{red}{?}}\)

OpenStudy (madeleine2):

3/2=1.5

OpenStudy (3mar):

I was busy. Sorry!

OpenStudy (3mar):

With @563blackghost , you are in good hands!

563blackghost (563blackghost):

Correct now we must determine the y-intercept. So we would plug this into slope-intercept form and plug one point into the equation to find `b`. \(\huge\bf{1=\frac{3}{2}(1)+b}\) Simplify to find b.

OpenStudy (3mar):

@madeleine2 Do you follow?

563blackghost (563blackghost):

`1 by 3/2 = 3/2` \(\huge\bf{1=\frac{3}{2}+b}\) Now subtract 1 and 3/2. \(\huge\bf{-\frac{1}{2}=b}\) So we plug this into slope-intercept form. \(\huge\bf{y=\frac{3}{2}x-\frac{1}{2}}\) We now would turn this into standard form. First we must rid of the fraction in order to do that we would multiply by the least common denominator. So in this case 2. \(\huge\bf{2y=2(\frac{3}{2}x-\frac{1}{2})}\) \(\huge\bf\color{orange}{2y=3x-1}\) Next subtract 3x from each side. What is your standard form equation? @madeleine2

563blackghost (563blackghost):

Can you please cooperate @madeleine2 ?

OpenStudy (3mar):

Nice presentation! Thank you for good explanation & illustration.

563blackghost (563blackghost):

Thank you 3mar :)

OpenStudy (3mar):

You are welcome, BEST SISTER & BEST TEACHER!